a solution of 36% of water and 64% acetaldehyde by mass.calculate the mole fraction of acetaldehyde

Grade:11

5 Answers

akash luckey aggara
11 years ago

8/19

 

Nishant Sharma
11 years ago

Let take the solution of 100 gram

Oxygen =36gram

Acetaldehyde = 64gram 

noxygen=36/18=2mole=n1

acetaldehyde=64/44=1.5=n2

Mole fraction of acetaldehyde= n2/(n1+n2)=0.41

TANAYRAJ SINGH CHOUHAN
11 years ago

Let us consider 100gm of solution

So the solution contains 36gm water and 64gm acetaldehyde(C2H4O)

So no. of moles of water = 36/18 = 2moles

& no. of moles of acetaldehyde = 64/44 = 16/11 moles

Now mole fraction of water = 2/(2+16/11) = 2*11/38 = 11/19

& mole fraction of acetaldehyde = 16*11/11*38 = 8/19

dattu savili
11 years ago

acetaldehyde mole fraction is 0.64

GOPICHAND BODA
11 years ago

mole fraction of CH3COOH is 0.347